OpenStudy (anonymous):

Need Help please. Brian is creating a collage on a piece of cardboard that has an area of 150r2 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of the square root of r cubed square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

OpenStudy (mathstudent55):

150r^2/(sqrt(r))^3

hartnn (hartnn):

Total area covered = Area of each piece * number of pieces here, total area is 150 r^2 and Area of each piece given is \(\sqrt r^3\)

OpenStudy (anonymous):

I don't understand what to do.

hartnn (hartnn):

did u get this ? Total area covered = Area of each piece * number of pieces

OpenStudy (anonymous):

Yes!

hartnn (hartnn):

good! now if you have read the question, we are given with total area is 150 r^2 and Area of each piece = square root of r cubed square right ? just plug these in the formula

OpenStudy (anonymous):

What is the formula?

hartnn (hartnn):

Total area covered = Area of each piece * number of pieces

OpenStudy (anonymous):

How do I multiply them correctly?

hartnn (hartnn):

we need total number of pieces Total area covered = Area of each piece * number of pieces gives Number of pieces = Total area covered / Area of each piece =\(\large \dfrac{150r^2}{\sqrt{r^3}}\) got this?

OpenStudy (anonymous):

Yes, but how do I multiply them?

hartnn (hartnn):

first simplify the denominator, \(\sqrt {r^3}= \sqrt {r^2}\times \sqrt r =....?\) can u simplify ?

OpenStudy (mathstudent55):